Demand function is Q=1100-40P. Determine the quantity that results in maximum revenue
1
Expert's answer
2016-02-17T00:00:57-0500
Q=1100-40P, 40P = 1100 - Q, P = 27.5 - Q/40 - inverse demand function.
Total revenue is maximized, when marginal revenue MR = 0.
MR = TR' = (P*Q)' = (27.5Q - Q^2/40)' = 27.5 - Q/20
MR = 0, so 27.5 - Q/20 = 0,
Q/20 = 27.5
Q = 550 units.